OpenStudy (anonymous):

Solve 56x + x^3 = 15x^2 by factoring

OpenStudy (anonymous):

Wouldn't it be x(56+x^2) = (I don't know the rest)

OpenStudy (anonymous):

i think i got it: x(56+x^2)=15x divide by x both sides so: x(56+x^2)/x=15x/x then you get 56+x^2=15x arrange:i.e x^2-15x+56=0 1*56=56 -8+-7=-15 insert:x^2-8x-7x+56=0 x is common.... x(x-8)......-7(x-8)=0 (x-7)...(x-8)=0 since (x-7)=0 x=7 since (x-8)=0 x=8 possible values of x are: 8 & 7

OpenStudy (anonymous):

sorry the division part is suppposed to be: x(56+x^2)=15x^2 then divide by x both sides you get: 56+x^2=15x the rest is ok